A garden hose has a water flow rate of 20.50 liters per minute. How long will it take to fill an 8 foot x 10 foot swimming pool to a depth of 2.50 feet? 1 liter = 0.26417 gal, and 1 cubic foot = 7.4805 gal.
1
Expert's answer
2019-09-17T04:55:30-0400

8 "\times" 10"\times" 2.50 = 200 cubic feet

1 cubic foot = 7.4805 gal so 200 will be 200"\times" 7.4805 = 1496.1 gal

1 liter = 0.26417 gal, therefore, 1496.1 gal will be "\frac{1496.1}{0.26417}" = 5663.3986 liters


20.50 liters per minute, 5663.3986 will take "\frac{5663.3986}{20.50}" = 276.2633 minutes